OpenStudy (anonymous):

A pair of fair dice is tossed. Given that the sum is 6, find the probability that one of the dice shows 2. (Hint: Possible pairs to achieve a 6, (1, 5), (2, 4), …)

OpenStudy (anonymous):

2/5

OpenStudy (anonymous):

may i know how did you get the answe ? thank you !

OpenStudy (anonymous):

P(A|B) = P(A U B)/ P(B) = 2/36 / 5 / 36 = 2/5

OpenStudy (anonymous):

got it??

OpenStudy (anonymous):

erm . how did you get 2/36 and 5/36 ?

OpenStudy (anonymous):

P (A U B) means the probability of getting a 2 on one of the dice when uve got a six as the sum now the cases are (2,4) and (4,2) so only 2 cases out of 36 and the cases of getting sum 6 are 5/36 so

OpenStudy (anonymous):

another way of doing it is

OpenStudy (anonymous):

how maqny points give a sum of 6?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

3 ways

OpenStudy (anonymous):

5 ways 1,5 5,1 3,3 2,4 4,2

OpenStudy (anonymous):

oh ya.

OpenStudy (anonymous):

so now out of these 5, how many cases r there in which theres a 2 on either die?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

so there u go 2/5

OpenStudy (anonymous):

okay thanks a lot !!! i get it !

OpenStudy (amistre64):

and as an aside: I believe its: P(A|B) = P(AnB)/P(B)